JoyLoveReynolds:

Drag the tiles to the correct boxes to complete the pairs. Not all tiles will be used. Match each quadratic equation with its solution set.

JoyLoveReynolds:

fredo:

2x2-32=0 x=-4,4

fredo:

4x2-100=0 x=5,-5

fredo:

x2-55=9 x=-8,8

JoyLoveReynolds:

do you know he last one

ThatCreepyBoy:

I'm figuring it out one second

ThatCreepyBoy:

@fredo Get your butt on here and help joy out T^T

fredo:

I thought you had it -_-

ThatCreepyBoy:

Well I don't so go help him!

JoyLoveReynolds:

im a girl......

ThatCreepyBoy:

Sorry v.v, her"

fredo:

x2-140=-19 x=11,-11

ThatCreepyBoy:

That wasn't so hard now was it?

JoyLoveReynolds:

thank you

fredo:

Not at all i just thought u were doing it lol

ThatCreepyBoy:

I got lost x,D, thank you for helping tho.